# StarGen: A Spatiotemporal Autoregression Framework with Video Diffusion Model for Scalable and Controllable Scene Generation

### v1

@@ Abstract
Figure 1: Overview of the proposed method: (a) We introduce a spatiotemporal autoregression framework for long-range scene generation. The generated scene is represented as a set of sparsely sampled posed images. The generation of the current sliding window of images (blue dotted box) is conditioned on spatially adjacent images that are generated previously (green frustums) and temporally overlapping image (blue solid box) from the preceding window. (b) The spatial conditioning images are processed by a large reconstruction model, which extracts the 3D information and renders the reconstructed latent features to each novel view. These spatial features, together with the temporal conditioning image, are used to condition the generation of the current window through a video diffusion model and a ControlNet. (c) The framework is used to implement three downstream tasks, including sparse view interpolation, perpetual view generation, and layout-conditioned city generation.
[…]
@@ 3 Method
Figure 2: Spatiotemporal-Conditioned Video Generation. Given two posed images as spatial conditions (green box on the left), the large reconstruction model regresses their depth maps and feature maps. The two feature maps $\mathbf{F}_{i_{1}}^{\text{spat}}$ and $\mathbf{F}_{i_{2}}^{\text{spat}}$ are rendered into novel views $\mathbf{F}^{\text{nov}}$ and temporally compressed to the latent space of CogVideoX, resulting in $\mathbf{z}^{\text{spat}}$. Simultaneously, the temporal conditioning image (blue box on the right) is encoded to replace the corresponding latent in $\mathbf{z}^{\text{spat}}$, resulting in the spatiotemporal condition $\mathbf{z}^{\text{st-cond}}$, which conditions the generation of CogVideoX through a ControlNet.
@@ 3.2 Spatiotemporal-Conditioned Video Generation
Spatial Condition. Inspired by latenSplat [60], we predict the scene structure from the two spatial conditioning images, and render the reconstructed latent features into each novel view to guide the generation. Specifically, we utilize a Large Reconstruction Model (LRM) to predict scene structure from two spatial conditioning images $({\mathbf{I}_{i_{1}}^{\text{spat}}},{\mathbf{I}_{i_{2}}^{\text{spat}}})$. Unlike existing methods [13, 56, 6], which rely solely on color information, our approach incorporates a prior depth map and Plücker coordinates [47] as additional inputs. The complete LRM input is $\{\mathbf{I}^{\text{spat}}_{i},\hat{\mathbf{D}}^{\text{spat}}_{i},\hat{\textbf{P}}^{\text{spat}}_{i}\}_{i=i_{1},i_{2}}$, including the RGB image $\mathbf{I}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 3}$, the depth map $\hat{\mathbf{D}}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 1}$ predicted by Depth Anything V2 [67], and the Plücker coordinates $\hat{\mathbf{P}}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 6}$ derived from the input pose ${\mathbf{P}_{i}}$, where $H$ and $W$ are image height and width respectively. Following GS-LRM [75], we stack, patchify, and concatenate the LRM inputs into a sequence of tokens. These tokens are fed into a transformer network to regress the depth maps $\{\mathbf{D}_{i}^{\text{spat}}\}_{i=i_{1},i_{2}}$ and feature maps $\{\mathbf{F}_{i}^{\text{spat}}\}_{i=i_{1},i_{2}}$ corresponding to the two conditioning views. Note that, unlike the scale-free $\hat{\mathbf{D}}_{i}^{\text{spat}}$ obtained from monocular depth prediction, the regressed $\mathbf{D}_{i}^{\text{spat}}$ is expected to be scale-aligned with the input poses. So we can render the features into novel views given the novel view poses $\mathbf{P}^{\text{nov}}=\{\mathbf{P}_{j}\}_{j=1}^{N}$, obtaining the novel view features $\mathbf{F}^{\text{nov}}=\{\mathbf{F}_{j}^{\text{nov}}\}_{j=1}^{N}$ and depth maps $\mathbf{D}^{\text{nov}}=\{\mathbf{D}_{j}^{\text{nov}}\}_{j=1}^{N}$:
@@
Spatiotemporal Condition. Previous methods typically support temporal conditioning by tuning the backbone T2V model into an I2V model [68], resulting in the need to train different backbones for conditioning at different positions [63]. To enhance the versatility without backbone tuning, we directly replace the corresponding latent frame in $\mathbf{z}^{\text{spat}}$. Specifically, the temporal conditioning image $\mathbf{I}_{k}^{\text{temp}}$ is fed into the pretrained VAE encoder to obtain a latent $\mathbf{z}_{k}^{\text{temp}}$. For conditioning on the first frame, the first latent frame in $\mathbf{z}^{\text{spat}}$ is directly replaced with $\mathbf{z}_{k}^{\text{temp}}$. For conditioning on other frame, we modify the novel view pose trajectory by inserting three duplicated poses at frame $k$. These four static frames will be compressed into one in $\mathbf{z}^{\text{spat}}$, which is replaced with $\mathbf{z}_{k}^{\text{temp}}$. The replaced latent is denoted as $\mathbf{z}^{\text{st-cond}}$, which serves as the condition for CogVideoX through a ControlNet [10] to generate a video consistent with the spatial and temporal conditions.
Training Loss. The training loss combines three weighted terms:
@@
$\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{\mathbf{x},\,t,\,\mathbf{T},\,\mathbf{z}^{\text{st-cond}},\,\epsilon}\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},t,\mathbf{T},\mathbf{z}^{\text{st-cond}}\right)-\epsilon\right\|^{2},$ (7)
@@ 3.3 Downstream Tasks
Sparse view interpolation is a key step in sparse view reconstruction, as demonstrated by concurrent works [40, 71] closely related to ours. Given the start and end frames, it generates the intermediate images. In the framework of StarGen, we distinguish two cases based on overlap between the two input images. In cases where the two input images have large overlapping regions, they serve as both spatial and temporal conditions, which are fed into the spatiotemporal-conditioned video generation model to generate an interpolated video. For long-range scenes where the start and end frames share minimal or even no common region, we propose a two-pass approach. In the first pass, the process is similar to the previous approach, but the pose difference between adjacent generated images is larger compared to the first case, resulting in a set of sparsely sampled images for the next pass. In the second pass, each pair of adjacent images from the first pass is treated as the start and end frames of a clip. All clips are then generated sequentially using the proposed spatiotemporal autoregression.
Perpetual view generation is the task of generating novel views of a scene from a single image while allowing pose control [39, 71]. In the framework of StarGen, the input image serves as both the spatial and temporal condition for the first generated clip. The remaining clips are then generated using the proposed spatiotemporal autoregression. Compared to the task of sparse view interpolation, where both the start and end frames provide constraints, perpetual view generation only has a constraint on the first frame. As a result, it is more sensitive to error accumulation and places higher demands on the scalability of the generation model, making it an excellent demonstration of StarGen’s capabilities.
@@ 4.1 Experiment Setup
Implementation Details. Our model is composed of a large reconstruction model (LRM), a causal compression network (CCN), and a video diffusion model (VDM) with ControlNet. For the LRM, we use vision transformer architecture [17], which consists of 12 layers with a hidden size of 768, an MLP size of 4096, and 12 attention heads, totaling 114M parameters. For the CCN, we implement a Conv3d layer with both input and output channels set to 16, a kernel size of (3, 3, 3), a stride of (1, 1, 1), a dilation of (1, 1, 1), and a padding of (0, 0, 0). For the VDM, we use the pretrained CogVideoX-2B-T2I [68] without any fine-tuning. The ControlNet model contains 6 layers, which are the trainable copies of the first 6 layers of CogVideoX-2B-T2I.
Training Details. For the efficiency of training, the proposed model is trained at a resolution of 256$\times$256. The training process includes: 1) training the LRM+CCN from scratch using consecutive frames with a batch size of 384, starting with 1K warm-up steps and continuing for a total of 40K steps; 2) training the LRM+CCN using $1\sim 3$ frame intervals, with a batch size of 384 for 20K steps; 3) joint training of the LRM+CCN and ControlNet, also using $1\sim 3$ frame intervals, with a batch size of 240 for 15K steps. For the task of layout-conditioned city generation, we additionally train two ControlNets for depth and semantic control, using a batch size of 128 over 20K steps. We also fine-tune our model by training the LRM+CCN with a batch size of 16 for 11K steps, followed by separately training the ControlNet with a batch size of 80 for another 11K steps.
@@ 4.5 Ablation Study
Spatiotemporal-Conditioned Video Generation. We conduct ablation experiments on the task sparse view interpolation with single-clip videos to evaluate the impact of each design choice in the proposed spatiotemporal-conditioned video generation, as shown in Tab. 4. “w/ DUSt3R” replaces the proposed LRM with DUSt3R, as done in ViewCrafter. “w/o spatial cond.” refers to a pure video interpolation method where only the input images are fixed, without the inclusion of LRM. “w/o depth input” excludes depth maps predicted by DepthAnything as input, while “w/o depth loss” omits the novel view depth loss. Finally, “fix LRM” involves freezing the LRM parameters in the third training stage, training only the ControlNet. The results indicate that excluding any of these design choices leads to performance degradation.

### v3

@@ Abstract
Figure 1: Overview of the proposed method: (a) We introduce a spatiotemporal autoregression framework for long-range scene generation. The generated scene is represented as a set of sparsely sampled posed images. The generation of the current sliding window of images (blue dotted box) is conditioned on spatially adjacent images (green frustums) and temporally overlapping image (blue solid box). (b) Spatial conditioning images are processed by a large reconstruction model, which extracts the 3D information and renders the reconstructed latent features to each novel view. These spatial features, together with the temporal conditioning image, are used to condition the generation of the current window through a video diffusion model and a ControlNet. (c) The framework is used to implement three downstream tasks, including sparse view interpolation, perpetual view generation, and layout-conditioned city generation.
indicates equal contribution.indicates corresponding author. This work was partially supported by the National Key Research and Development Program of China (No. 2023YFF0905104) and NSF of China (No. 62425209).
@@ 3 Method
Figure 2: Spatiotemporal-Conditioned Video Generation. Given two posed images as spatial conditions (green dotted box on the left), the large reconstruction model regresses their depth maps and feature maps. The two feature maps $\mathbf{F}_{i_{1}}^{\text{spat}}$ and $\mathbf{F}_{i_{2}}^{\text{spat}}$ are rendered into novel views $\mathbf{F}^{\text{nov}}$ and temporally compressed to the latent space of CogVideoX, resulting in $\mathbf{z}^{\text{spat}}$. Simultaneously, the temporal conditioning image (blue dotted box on the right) is encoded to $\mathbf{z}^{\text{temp}}_{k}$ to replace the corresponding latent in $\mathbf{z}^{\text{spat}}$, resulting in the spatiotemporal condition $\mathbf{z}^{\text{st}}$, which conditions the generation of CogVideoX through a ControlNet.
@@ 3.2 Spatiotemporal-Conditioned Video Generation
Spatial Condition. Inspired by latenSplat [60], we predict the scene structure from the two spatial conditioning images, and render the reconstructed latent features for each novel view to guide the generation. Specifically, we utilize a Large Reconstruction Model (LRM) to predict scene structure from two spatial conditioning images $({\mathbf{I}_{i_{1}}^{\text{spat}}},{\mathbf{I}_{i_{2}}^{\text{spat}}})$. Unlike existing methods [13, 56, 6], which rely solely on color information, our approach incorporates a prior depth map and Plücker coordinates [47] as additional inputs. The complete LRM input is $\{\mathbf{I}^{\text{spat}}_{i},\hat{\mathbf{D}}^{\text{spat}}_{i},\hat{\textbf{P}}^{\text{spat}}_{i}\}_{i=i_{1},i_{2}}$, including the RGB image $\mathbf{I}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 3}$, the depth map $\hat{\mathbf{D}}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 1}$ predicted by Depth Anything V2 [67], and the Plücker coordinates $\hat{\mathbf{P}}^{\text{spat}}_{i}\in\mathbb{R}^{H\times W\times 6}$ derived from the input pose ${\mathbf{P}_{i}}$, where $H$ and $W$ are image height and width respectively. Following GS-LRM [75], we stack, patchify, and concatenate the LRM inputs into a sequence of tokens. These tokens are fed into a transformer network to regress the depth maps $\{\mathbf{D}_{i}^{\text{spat}}\}_{i=i_{1},i_{2}}$ and feature maps $\{\mathbf{F}_{i}^{\text{spat}}\}_{i=i_{1},i_{2}}$ corresponding to the two conditioning views. Note that, unlike the scale-free $\hat{\mathbf{D}}_{i}^{\text{spat}}$ obtained from monocular depth prediction, the regressed $\mathbf{D}_{i}^{\text{spat}}$ is expected to be scale-aligned with the input poses. So we can render the features into novel views given the novel view poses $\mathbf{P}^{\text{nov}}=\{\mathbf{P}_{j}\}_{j=1}^{N}$, obtaining the novel view features $\mathbf{F}^{\text{nov}}=\{\mathbf{F}_{j}^{\text{nov}}\}_{j=1}^{N}$ and depth maps $\mathbf{D}^{\text{nov}}=\{\mathbf{D}_{j}^{\text{nov}}\}_{j=1}^{N}$:
@@
Spatiotemporal Condition. Previous methods typically support temporal conditioning by tuning the backbone T2V model into an I2V model [68], resulting in the need to train different backbones for conditioning at different positions [63]. To enhance the versatility without backbone tuning, we directly replace the corresponding latent frame in $\mathbf{z}^{\text{spat}}$. Specifically, the temporal conditioning image $\mathbf{I}_{k}^{\text{temp}}$ is fed into the pretrained VAE encoder to obtain a latent $\mathbf{z}_{k}^{\text{temp}}$. For conditioning on the first frame, the first latent frame in $\mathbf{z}^{\text{spat}}$ is directly replaced with $\mathbf{z}_{k}^{\text{temp}}$. For conditioning on other frame, we modify the novel view pose trajectory by inserting three duplicated poses at frame $k$. These four static frames will be compressed into one in $\mathbf{z}^{\text{spat}}$, which is replaced with $\mathbf{z}_{k}^{\text{temp}}$. The replaced latent is denoted as $\mathbf{z}^{\text{st}}$, which serves as the condition for CogVideoX through a ControlNet [11] to generate a video consistent with the spatial and temporal conditions.
Training Loss. The training loss combines three terms:
@@
$\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{\mathbf{x},\,t,\,\mathbf{T},\,\mathbf{z}^{\text{st}},\,\epsilon}\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},t,\mathbf{T},\mathbf{z}^{\text{st}}\right)-\epsilon\right\|^{2},$ (7)
@@ 3.3 Downstream Tasks
Sparse view interpolation is a key step in sparse view reconstruction, as demonstrated by concurrent works [40, 71] closely related to ours. Given the start and end frames, it generates the intermediate images. In the framework of StarGen, we distinguish two cases based on overlap between the two input images. In cases where the two input images have large overlapping regions, they serve as both spatial and temporal conditions, which are fed into the spatiotemporal-conditioned video generation model to generate an interpolated video. For long-range scenes where the start and end frames share minimal or even no common region, we propose a two-pass approach. In the first pass, the process is similar to the previous approach, but the pose difference between adjacent generated images is larger compared to the first case, resulting in a set of sparsely sampled images. In the second pass, each pair of adjacent images from the first pass is treated as the start and end frames of a clip. All clips are generated as the first case.
Perpetual view generation is the task of generating novel views of a scene from a single image while allowing pose control [39, 71]. In the framework of StarGen, the input image serves as both the spatial and temporal condition for the first generated clip. The remaining clips are then generated using the proposed spatiotemporal autoregression. Compared to the task of sparse view interpolation, where both the start and end frames provide constraints, perpetual view generation only has a constraint on the first frame. As a result, it is sensitive to error accumulation and places high demands on the scalability of the generation model.
@@ 4.1 Experiment Setup
Implementation Details. Our model is composed of a large reconstruction model (LRM), a causal compression network (CCN), and a video diffusion model (VDM) with ControlNet. For the LRM, we use vision transformer architecture [17], which consists of 12 layers with a hidden size of 768, an MLP size of 4096, and 12 attention heads, totaling 114M parameters. For the CCN, we implement a Conv3d layer with both input and output channels set to 16, a kernel size of (3, 3, 3), a stride of (1, 1, 1), a dilation of (1, 1, 1), and a padding of (0, 0, 0). For the VDM, we use the pretrained CogVideoX-2B [68] without any fine-tuning. The ControlNet model contains 6 layers, which are the trainable copies of the first 6 layers of CogVideoX-2B. Note that different baseline methods employ different VDMs, which may introduce unfair comparisons. To mitigate this flaw, we provide an ablation study in the supplementary material (Sec. 6) to demonstrate that the improvement over the baseline methods is not solely attributable to the stronger VDM backbone.
Training Details. For the efficiency of training, the proposed model is trained at a resolution of 256$\times$256. The training process includes: 1) training the LRM+CCN from scratch using consecutive frames with batch size 384, starting with 1K warm-up steps and continuing for a total of 40K steps; 2) training the LRM+CCN using $1\sim 3$ frame intervals, with batch size 384 for 20K steps; 3) joint training of the LRM+CCN and ControlNet, also using $1\sim 3$ frame intervals, with batch size 240 for 15K steps. For the task of layout-conditioned city generation, we additionally train two ControlNets for depth and semantic control, with batch size 128 over 20K steps. We also fine-tune our model by training the LRM+CCN with batch size 16 for 11K steps, followed by separately training the ControlNet with a batch size of 80 for another 11K steps.
@@ 4.5 Ablation Study
Spatiotemporal-Conditioned Video Generation. We conduct ablation experiments on the task sparse view interpolation with single-clip videos to evaluate the impact of each design choice in the proposed spatiotemporal-conditioned video generation, as shown in Tab. 4. “w/ DUSt3R” replaces the proposed LRM with DUSt3R, as done in ViewCrafter. “w/o spatial cond.” refers to a pure video interpolation method where only the input images are fixed, without the inclusion of LRM. “w/o depth input” excludes depth maps predicted by Depth Anything V2 as input, while “w/o depth loss” omits the novel view depth loss. Finally, “fix LRM” involves freezing the LRM parameters in the third training stage, training only the ControlNet. The results indicate that excluding any of these design choices leads to performance degradation.